# SimpleToM: Exposing the Gap between Explicit ToM Inference and Implicit ToM Application in LLMs

### v1

@@ 5.3 No applied ToM in LLMs? Exploring the rabbit hole of human hand-holding
3. Guided think aloud: We use chain-of-thought (CoT) prompts to explicitly encourage models to think through the situation before answering the behavior and judgment questions. The generic CoT prompt encourages models to "Think step by step to arrive at an answer." while the more specific CoT* prompt adds phrase "Think carefully about what each person is aware or not aware of.". [^9] See detailed prompts in Appendix G.3. The results are shown in Table 5. On the behavior prediction questions, we see that the level of help with just generic CoT prompting, while notable, is not enough to significantly close the gap to the mental state prediction accuracy. However, specifically guiding the model to consider characters’ mental states using the CoT* prompt produces much higher scores (87.4% to 92.7% accuracy across the models). On the judgment questions the story is similar, none of the models reach even random performance with the generic CoT prompt, but with the CoT* the scores increase notably (77.8% to 86.7%) while still remaining significantly below the mental state scores.
4. Combine reminder and thinking aloud: We finally combine the two most effective interventions by including the mental state reminder with the CoT* chain-of-thought prompt, also recorded in Table 5. With this combined approach, all three models produce high scores across the board for both the behavior and judgment questions. In fact the Claude-3.5-Sonnet model reaches an average score of 97.1% with this method, serving as a quality check of SimpleToM, since with enough reminders and (seemingly obvious) hints, near-perfect scores are achieved.
We include examples of chain-of-thought outputs in Appendix I, illustrating how the reasoning can go wrong when an insufficient level of intervention is provided. Figure 9 shows how GPT-4o with generic CoT has the faulty reasoning “Given that the toolbox contains a dangerous snake, the neighbor’s primary concern would likely be safety”, without considering percepts and mental state. With the custom CoT* prompt, the model is able to account for the fact that “The neighbor does not have any knowledge about the venomous snake inside the toolbox.” Figure 10 shows that if not explicitly reminded of the mental state question, Claude can erroneously conclude that the “correct” behavior can be judged as unreasonable “regardless of the awareness of the specific issue.”
6 Related Work
Theory of Mind has been extensively studied in psychology in a range of scenarios (see Appendix  B). ToM reasoning, and broadly social commonsense, has also been shown to be important by the different parts of the AI community including in conversations (Kim et al., 2023b; a), games (Zhou et al., 2023b), and even multi-modal setups (Jin et al., 2024), with most popular ToM tests using stories to probe LLMs. Relying on stories from small test sets in cognitive science studies to benchmark ToM abilities in LLMs (Bubeck et al., 2023; Kosinski, 2024) could produce results that differ given minor alterations (Ullman, 2023) and would be more robust if tested on larger samples. Yet expert-crafted or naturally occurring self-contained stories that can serve as targeted tests of ToM are scarce and human story-writing is expensive, leading to the use of automatically generated datasets for studying ToM behavior in LLMs (Jung et al., 2024; Wilf et al., 2023; Sap et al., 2023; Shapira et al., 2023; Sclar et al., 2023). Existing generated datasets allow studies of ToM to be carried out at scale, but templated stories often limit settings where information asymmetry arises. For example, the entire dataset might only contain stories about some object being moved (over-reliance on classical Sally-Anne task, e.g., in ToMi (Le et al., 2019), ToM-bAbI (Nematzadeh et al., 2018), Hi-ToM (Wu et al., 2023), OpenToM (Xu et al., 2024)). Or the dataset might focus on whether some character has witnessed a sudden external event (BigToM (Gandhi et al., 2024)). These stories are often systematically generated with the explicit use of mentalizing words to convey percepts and beliefs, e.g., “unknown to Amy” and “Amy thinks that” in Xu et al. (2024) or “Noor sees” and “Mei does not notice” in Gandhi et al. (2024). However, the explicit use of mentalizing words also makes the stories (i) unnaturally simplistic, having removed the need for commonsense inferences about percepts or beliefs, and (ii) sometimes unrealistic, with combinations like “Cheng does not notice the power outage” when he “use[s] a projector to show a documentary”(Gandhi et al., 2024). Other existing datasets could be improved by addressing issues such as lacking exploration of applied ToM beyond action prediction (Zhou et al., 2023a; Gandhi et al., 2024), confounding factors like memory loads or tracking requirements (Le et al., 2019; Xu et al., 2024), and violating Quesque & Rossetti (2020)’s criteria (see Appendix  B) for validating ToM (Chen et al., 2024). Our work extends existing datasets by following Tian et al. (2024) in combining the generative strength of LLMs and the verification ability of human annotators, and extends the existing efforts toward robust, generalizable evaluation (Kiela et al., 2021; Srivastava et al., 2024), avoiding known pitfalls while preserving the systematic and scalable nature of the dataset creation process.
7 Conclusion
SimpleToM is the first dataset of its kind testing both explicit and applied ToM using a large set of diverse, concise, simple stories, covering different ways in which information asymmetry may arise. The dataset opens up new opportunities for evaluating and debugging ToM abilities of LLMs, especially in the relatively under-studied area of applied ToM. Our analyses reveal a jarring gap between explicit and applied ToM capabilities in current frontier LLMs. Thus, if our goal is LLM agents capable of applying ToM in complex, human-centered environments, we need to look beyond testing LLMs with psychology-inspired ToM questions, and also start testing them more rigorously on applied ToM (e.g., behavioral prediction and judgment) in different situations.
We show that this gap can be largely closed in the best models by well-designed interventions at inference time, like reminding the model of its answers to important questions or guiding it with a custom chain-of-thought prompt. We argue that a robust LLM should perform well on SimpleToM without such interventions, so it can independently and flexibly apply ToM-related reasoning whenever required within potentially complex and multi-faceted environments. Model developers interested in real-world deployment of their models, should be alert to closing this performance gap so the models can interact with society appropriately, ideally without the higher inference costs associated with explicit chain-of-thought reasoning or implicit o1-preview reasoning tokens (see further discussion in Appendix H).
SimpleToM can also facilitate the community in pursuing various exciting directions for future work, including studying how ToM performance may differ with stories that involve different levels of harmfulness and unethicality (see Appendix E.3), and innovative modeling approaches that can help close the gap between explicit and applied ToM in AI models.
Ethics Statement
Reproducibility
We make our SimpleToM dataset and the full evaluation data for the analyzed models publicly available. This will allow researchers to reproduce and build on top of our work in studying the neural ToM capabilities of LLMs.
Further, we provide all prompts used for SimpleToM creation – see Appendix E.5 for the entity brainstorming prompt, and Appendix E.4 for the story generation prompt. We also carefully document the instructions used in our crowdsourcing process (Appendix D.1) and how we qualified workers (Appendix D.2). All prompts used for the different inference interventions are provided in Appendix G.
Appendix A FAQs
  • Q:
    How is SimpleToM different from existing datasets?
  • SimpleToM addresses limitations in previous efforts to examine Theory-of-Mind (ToM) reasoning in LLMs, by (1) having diverse false belief setups (e.g., beyond those in Sally-Ann task where some object is moved when a character is not present), (2) requiring LLMs to make commonsense inferences in situations rather explicit use of mentalizing words to convey what characters perceive or believe, and (3) going beyond explicit ToM to test models’ ability to apply inferred knowledge in follow-up applied ToM questions (such as behavior prediction and judgment of behavior).
  • Q:
  • Q:
    What do the different inference interventions reveal about models’ ToM capabilities?
  • Patch, guide, think aloud combined method results show belief-to-judgment inference robustness. When reminded of their answer to mental state questions about information awareness and using ToM-specific CoT, models’ belief-to-judgment inference is pretty robust in existing LLMs. The models have some knowledge linking irrationality and unawareness (e.g., unawareness should be accounted for seemingly irrational behavior). So arriving at the “correct belief” will be crucial part of ToM-based reasoning, but this is not done implicitly in current frontier models’ reasoning.
  • Q:
  • We illustrate in Appendix F some prompt variations that we have experimented with for the judgment question. Across Llama-3.1-405B, Claude-3.5-Sonnet and GPT-4o, the scores using different variants were all consistently below random (never exceeding 30% accuracy), indicating that the low scores on the judgment questions come more from fundamental flaws in the applied ToM capabilities of models rather than an effect of specific formatting/wording.
Appendix B Studies of ToM in psychology
Theory of Mind has been extensively studied in psychology in a range of scenarios, for instance, studies of manipulation, secrecy (Peskin & Ardino, 2003), deception, lying (Lewis et al., 1989; Perner, 1993; Peskin, 1992), misleading behavior (Chandler et al., 1989; Wimmer & Perner, 1983; Doherty, 2008), autism (Frith & Happé, 1994), and analysis of rational behavior (Gergely & Csibra, 2003; Liu & Spelke, 2017). Classical tests of ToM in developmental psychology include testing the development of this ability in children via false belief prediction – using the unexpected transfer false belief task, the Sally-Anne task (Baron-Cohen et al., 1985), or the unexpected contents false belief task, the Smarties task (Perner et al., 1987). Quesque & Rossetti (2020) review classic tests of ToM and outline two important criteria for tasks that validate ToM: (1) The task must indicate that the respondents can differentiate between the other’s mental state and their own. (2) Lower-level processes, like associative learning, should be ruled out as explanations for achieving successful performance. Given the wide applicability of ToM reasoning in various real-life situations such as analyzing people’s behavior (Liu et al., 2024; Jara-Ettinger et al., 2016) and making judgments (Schein & Gray, 2018; Young et al., 2007), there has also been increasing interest in assessing ToM capabilities in AI models (Le et al., 2019; Ullman, 2023; Kosinski, 2024; Jin et al., 2024; Trott et al., 2023).
Appendix C Details of LLMs used in experiments
Model Full name Provider
Llama-3.1-8B Meta-Llama-3.1-8B-Instruct-Turbo Meta
Claude-3-Haiku claude-3-haiku-20240307 Anthropic
GPT-3.5 gpt-3.5-turbo-1106 OpenAI
Claude-3-Opus claude-3-opus-20240229 Anthropic
GPT-4 gpt-4-0125-preview OpenAI
Llama-3.1-405B Meta-Llama-3.1-405B-Instruct-Turbo Meta
Claude-3.5-Sonnet claude-3-5-sonnet-20240620 Anthropic
GPT-4o gpt-4o-2024-05-13 OpenAI
Partially evaluated:
o1-mini o1-mini-2024-09-12 OpenAI
o1-preview o1-preview-2024-09-12 OpenAI
Table 6: Details of models used for evaluation and dataset creation. The very recent o1 models were partially evaluated under different conditions (unable to set temperature to zero).
Appendix D Details on crowdsourcing to ensure validity of stories for testing ToM
D.1 Instructions to crowdworkers
@@
The crowdsourcing instructions included a detailed description of the motivation behind the annotation task and what is to be annotated (see Figure 5). We also provide four detailed examples (Figures 6 and 7) for each of the aspects to annotate, illustrating and giving justifications for circumstances under which different annotation options would be appropriate. The workers were then asked to provide their own set of annotations when presented with story (and likely actions) using the question templates shown in Figure 8.
D.2 Qualification round
@@
D.3 Crowdworkers and pay rate
Our participants were recruited on the Amazon Mechanical Turk (AMT) platform. The workers that worked on our annotation task met minimum qualification in AMT of >=>=> =98% approval rate, with at least 10k approved HITs. They were from US locations and rated at Amazon’s Masters Level. They must also not have the record of having accepted but not complete a HIT posted by our AMT account. In addition to these qualifications, participants of the actual annotation task (on the 3600 generated stories) must have also passed our rigorous qualification task described above (Appendix D.2). The workers were paid at a rate of $\approx$\$15/hr.
@@ D.3 Crowdworkers and pay rate
D.4 Strict quality filter
@@
Appendix E Details of SimpleToM
E.1 Scenarios in SimpleToM
@@
Table 7: Description and examples for broad scenarios where information asymmetry occurs naturally in the real world.
E.2 Entities in SimpleToM
@@ E.3 Harmfulness and unethicality
E.4 Story generation prompt
@@
E.5 Entity brainstorming prompt
@@
Appendix F Discussion on prompt variations
@@
Appendix G Details of different inference interventions
We illustrate the interventions used to give LLMs various levels of hints, such as reminding the models about their mental state answers, using chain-of-thought prompts to encourage careful reasoning, or doing so via the system prompt
G.1 Patching mental state inference in the prompt (MS remind)
@@
G.2 Guiding via system prompt (SysP)
@@
G.3 Encouraging reasoning/thinking before answering (CoT)
@@
Appendix H Inference costs from interventions
@@
Appendix I Examples of chain-of-thought outputs
@@ 6. Analysis:
Appendix J Performance across scenarios
@@
Figure 13: Analyzing where top models fail first in the the sequence of predicting mental state, then behavior and finally judgment (see Figure 3 for legend). We can record failures for the first mistake e.g., whether models (i) fail at the mental state (MS) question, (ii) pass the MS question but fail at behavior prediction, (iii) pass both MS and behavior questions but fail at judgment question, or (iv) get all 3 questions correct. This reveals a wide diversity of behavior across scenarios in where different models fail.
Is o1-preview always near perfect? Figure 4 shows how models differ in performance across different scenarios and question types. [^10] Figures 11 and  12 provide results across all broad scenarios for different models. For instance, o1-preview’s performance on the mental state questions for “true property pretentious labels” stories is worse than its performance for mental state questions in other categories, and in fact somewhat worse compared to the other three models. Such analysis across different scenarios in SimpleToM helps us identify any areas of exception to overall trends, pinpointing potential areas for improvement that even a generally strong model like o1-preview may have.
@@
Perfection is possible but many LLMs are not there. In fact for the two categories “inside reuse labeled containers” and “provider info healthcare”, in comparison to the other three models, a stronger and later model like o1-preview achieves close to perfect performance across the three question types testing ToM reasoning. This further illustrates the high-quality nature of SimpleToM, in that these simple two-sentence stories are clean, straightforward tests of neural ToM reasoning, yet models other than the strongest and latest model o1-preview shows poor performance on applied ToM questions (behavior and judgment) in various ways. Model developers, if interested in real-world deployment of their models, should be alert into closing this performance gap so as to ensure their models can interact with society appropriately, ideally without the high inference costs of chain-of-thought reasoning and o1-preview reasoning tokens (that will be discussed in Section 7 and Appendix H).

### v2

@@ Appendix F No applied ToM in LLMs? Exploring the rabbit hole of human hand-holding
Appendix G Details on crowdsourcing to ensure validity of stories for testing ToM
G.1 Instructions to crowdworkers
@@
The crowdsourcing instructions included a detailed description of the motivation behind the annotation task and what is to be annotated (see Figure 6). We also provide four detailed examples (Figures 7 and 8) for each of the aspects to annotate, illustrating and giving justifications for circumstances under which different annotation options would be appropriate. The workers were then asked to provide their own set of annotations when presented with story (and likely actions) using the question templates shown in Figure 9.
G.2 Qualification round
@@
G.3 Crowdworkers and pay rate
Our participants were recruited on the Amazon Mechanical Turk (AMT) platform. The workers that worked on our annotation task met minimum qualification in AMT of >=>=98% approval rate, with at least 10k approved HITs. They were from US locations and rated at Amazon’s Masters Level. They must also not have the record of having accepted but not complete a HIT posted by our AMT account. In addition to these qualifications, participants of the actual annotation task (on the 3600 generated stories) must have also passed our rigorous qualification task described above (Appendix G.2). The workers were paid at a rate of $\approx$\$15/hr.
@@ G.3 Crowdworkers and pay rate
G.4 Strict quality filter
@@
Appendix H Human baseline
Figure 10: Humans demonstrate greater consistency across the question types in SimpleToM. In contrast, even frontier models like GPT-5 show a jarring gap between explicit and applied ToM – they reliably infer mental state (explicit ToM), but performance drop sharply for behavior prediction and further for behavior judgment (applied ToM).
We sample 50 stories from SimpleToM, covering 5 stories from each of the 10 broad scenario categories described in Section 2.1. Each story comes with 3 questions, testing mental state inference, behavior prediction and judgment. For each of the total of 150 questions, we gathered a human response using crowdsourcing (with a similar process as described in Appendix G but this time to estimate human performance on these sampled stories and questions).
Figure 10 shows that the human baseline is relatively consistent, with people maintaining strong accuracy when inferring mental states, predicting subsequent behavior, and making related judgments (significantly less performance drop compared to LLMs). In stark contrast, even state-of-the-art LLMs like GPT-5 exhibit a striking performance gap, excelling at the mental state inference task (explicit ToM) but failing to apply that knowledge with accuracy dropping sharply for applied ToM questions (behavior prediction and judgment). This salient gap between explicit ToM inference and implicit ToM application is observed across different LLMs in Table 2. Our results reveal that while language models can explicitly infer mental states, they cannot reliably apply that inference and reason about the implications in the same consistent way humans do.
Appendix I Details of SimpleToM
I.1 Scenarios in SimpleToM
@@
Table 8: Description and examples for broad scenarios where information asymmetry occurs naturally in everyday scenarios.
@@
I.2 Entities in SimpleToM
@@ I.3 Harmfulness and unethicality
I.4 Story generation prompt
@@
I.5 Entity brainstorming prompt
@@
Appendix J Discussion on prompt variations
@@
Appendix K Details of different inference interventions
We illustrate the interventions used to give LLMs various levels of hints, such as reminding the models about their mental state answers, using chain-of-thought prompts to encourage careful reasoning, or doing so via the system prompt.
K.1 Patching mental state inference in the prompt (MS remind)
@@
K.2 Guiding via system prompt (SysP)
@@
K.3 Encouraging reasoning/thinking before answering (CoT)
@@
Appendix L Inference costs from interventions
@@
Appendix M Examples of chain-of-thought outputs
@@ 6. Analysis:
Appendix N Further analysis on interesting behavior
Llama-3.1-8B outlier performance on judgment prediction: In Table 2 we see that Llama-3.1-8B’s score of 54.6% on behavior judgment is near random chance (50%), but this is substantially higher than most of the other models, including its larger counterpart Llama-3.1-405B. This reveals the following insights about the Llama-3.1-8B model:
(1) It has less bias to being consistently wrong in the judgment task than the other models.
(2) Comparing the performance on behavior prediction and behavior judgment in more detail, there is much inconsistency within the behavior-to-judgment reasoning chain. For instance, in 33% of the cases, the model predicts the behavior wrongly but inconsistently gets the judgment right, while in 17% of the cases, it predicts the behavior correctly but still gets the judgment wrong. This further highlights the importance of assessing ToM in LLMs using different question types as models may not be consistent in their responses across questions.
o1-preview’s built-in inference-time reasoning tokens help with applied ToM: The built-in inference-time reasoning tokens are akin to the chain-of-thought responses, although lengthier, suggesting that the model is iterating on its reasoning towards a final answer. As noted in Appendix L, empirically we notice that the o1-preview model uses a lot more tokens than other models with CoT*. One hypothesis regarding o1-preview’s built-in inference-time reasoning tokens being helpful in applied ToM reasoning is that they go through a longer reasoning process, which could potentially involve backtracking or self-questioning along the way (mimicking human intervention), leading to somewhat better performance.
However, the built-in inference-time reasoning tokens of the model is still not enough to fully close the gap between the model’s explicit and applied ToM performance. This further highlights the novelty of the gap our paper exposes - even this recently released model, using a relatively large number of reasoning tokens to reason about simple 2-sentence stories, still shows a significant gap in explicit and applied ToM performance (see Table 2).
Appendix O Performance across scenarios
In Figures 13 and  14, we show how model performance varies across scenarios.
@@
Figure 15: Analyzing where top models fail first in the sequence of predicting mental state, then behavior and finally judgment (see Figure 3 for legend). We can record failures for the first mistake e.g., whether models (i) fail at the mental state (MS) question, (ii) pass the MS question but fail at behavior prediction, (iii) pass both MS and behavior questions but fail at judgment question, or (iv) get all 3 questions correct. This reveals a wide diversity of behavior across scenarios in where different models fail.
Is o1-preview always near perfect? Figures 13 and  14 show how models differ in performance across different scenarios and question types. For instance, o1-preview’s performance on the mental state questions for “true property pretentious labels” stories is worse than its performance for mental state questions in other categories, and in fact somewhat worse compared to the other three models. Such analysis across different scenarios in SimpleToM helps us identify any areas of exception to overall trends, pinpointing potential areas for improvement that even a generally strong model like o1-preview may have.
@@
Perfection is possible but many LLMs are not there. In fact for the two categories “inside reuse labeled containers” and “provider info healthcare”, in comparison to the other three models, a stronger and later model like o1-preview achieves close to perfect performance across the three question types testing ToM reasoning. This further illustrates the high-quality nature of SimpleToM, in that these simple two-sentence stories are clean, straightforward tests of neural ToM reasoning, yet models other than the stronger and later model o1-preview show poor performance on applied ToM questions (behavior and judgment) in various ways. Model developers, if interested in real-world deployment of their models, should be alert into closing this performance gap so as to ensure their models can interact with society appropriately, ideally without the high inference costs of chain-of-thought reasoning and o1-preview reasoning tokens (further discussed in Section 7 and Appendix L).
Appendix P Performance across personas
Another inference-time intervention is to imbue certain personas onto the models. We experiment with two of the models (GPT-4o and Claude-3.5-Sonnet), using the following five personality descriptions from PersonaHub: [^13] https://huggingface.co/datasets/proj-persona/PersonaHub
  • lawyer: You are a partner at the law firm, recognized for their extensive knowledge of healthcare laws.
  • worker: You are a factory worker who doesn’t trust the COVID-19 vaccine.
  • atheist: You are an atheist, philosophy lecturer who encourages open dialog about faith and belief systems.
  • psychology student: You are a university psychology student who is currently studying creativity and personality.
  • psychologist: You are a clinical psychologist collaborating with the music therapist to provide holistic patient care.
We inserted these into the system prompt to produce the results shown in Table 10. Across personas, we see that the gap between explicit ToM (mental state prediction) and applied ToM (behavior prediction and judgment) remains consistently prominent, similar to the case where no persona is specified. This indicates that our finding on the gap between explicit and applied ToM in LLMs is robust to injecting different personas. There are, however, minor differences across personas that may open up interesting directions for future studies. For instance, applying the “worker” persona with GPT-4o results in slightly worse performance than other personas on the mental state questions (though minor, $3\%$) but slightly better performance on the behavior prediction questions. It is also consistent across GPT-4o and Claude 3.5 (and more prominent in the case of Claude 3.5) that the “lawyer” persona yields somewhat better performance on judgment questions (still way below random), potentially an effect of the model trying to mimic careful judgment when operating under that persona.
Table 10: Evaluation results for SimpleToM on the different question types across 5 alternate personas, showing minor differences in scores, but without significantly closing the gap in performance between the explicit ToM mental state questions vs the implicit ToM behavior and judgment questions.
model persona mental state behavior judgment
(Explicit ToM) (Applied ToM) (Applied ToM)
GPT4-o 95.6 49.5 15.3
GPT4-o lawyer 95.5 49.7 17.2
GPT4-o worker 93.1 55.8 15.9
GPT4-o atheist 95.0 50.9 15.6
GPT4-o psychology student 94.4 47.6 15.5
GPT4-o psychologist 95.2 53.9 16.7
Claude-3.5-Sonnet 97.9 67.0 24.9
Claude-3.5-Sonnet lawyer 98.4 67.6 32.0
Claude-3.5-Sonnet worker 97.9 67.0 24.5
Claude-3.5-Sonnet atheist 97.9 65.8 23.4
Claude-3.5-Sonnet psychology student 97.3 68.6 24.8
Claude-3.5-Sonnet psychologist 97.9 68.6 26.9
Appendix Q Statistical details
Q.1 95% confidence intervals
In this section we report further statistical details for on our model evaluations. For accuracy values in Table 3 in the main text, we annotated with a 95% confidence interval in Table 11.
Q.2 Paired tests/bootstraps
Table 12 reports paired 95% bootstrap confidence intervals for differences between tasks. Because each model answers the same set of stories across tasks, we use paired bootstrapping: for each of 10,000 replicates we resample stories with replacement, recompute the corresponding task differences (e.g., MS–BP), and take the 2.5 and 97.5 percentiles of these bootstrap differences as the confidence interval. Below each interval we report a one-sided bootstrap $p$-value testing whether the gap is strictly positive (i.e., whether A–B >0>0). When fewer than 0.1% of bootstrap replicates reverse the sign of the difference, we report $p<0.001$; otherwise we show the empirical value. This quantifies how reliably each model exhibits the observed ToM gaps.
Across all models in Table 12, the bootstrap confidence intervals show a large and consistently positive MS–BP gap: models are substantially better at explicitly identifying mental states than at predicting agents’ behavior. For many models the MS–BP interval is wide (often 30–50 points), and $p<0.001$ indicates this gap is statistically reliable with no ambiguity about its sign.
The BP–JU column shows how much behavior prediction outperforms (positive intervals) or underperforms (negative intervals) judgment of behavior. Large negative intervals (e.g., in GPT-3.5 or Llama-3.1-8B) indicate that these models judge behaviors more accurately than they can predict them. Large positive intervals (e.g., in Claude models) indicate the reverse: prediction is easier for them than judging behavior.
The MS–JU column combines both gaps. Its very large positive CIs (often 60–90 points) show that no model comes close to carrying its explicit ToM ability through to judgment tasks.
The BP–0.5 and JU–0.5 columns test whether BP or JU are above or below chance. Intervals fully below zero (with $p<0.001$) show that many models perform significantly below chance, especially on JU. Intervals straddling zero (e.g., GPT-4o BP) show performance not distinguishable from random choice.
Overall, the table demonstrates that the explicit vs applied ToM gap is large, consistent across models, and statistically unambiguous, while many models’ applied ToM performance is at or even significantly below chance.
Table 11: Evaluation with guidance via mental state reminder (MS remind), system prompt guiding (SysP), and chain-of-thought prompting (CoT). Each score is annotated with a 95% confidence interval.
model MS behavior prediction judgment of behavior
intervention none none MS remind SysP CoT none MS remind SysP CoT
GPT-4o 95.6 ($\pm$1.2) 49.5 ($\pm$2.9) 82.8 ($\pm$2.2) 47.3 ($\pm$2.9) 62.8 ($\pm$2.8) 15.3 ($\pm$2.1) 42.2 ($\pm$2.9) 14.9 ($\pm$2.1) 39.2 ($\pm$2.8)
Llama-3.1-405B 97.8 ($\pm$0.8) 58.2 ($\pm$2.9) 89.5 ($\pm$1.8) 64.5 ($\pm$2.8) 57.2 ($\pm$2.9) 10.0 ($\pm$1.7) 25.8 ($\pm$2.5) 9.9 ($\pm$1.7) 35.2 ($\pm$2.8)
Claude-3.5-Sonnet 97.9 ($\pm$0.8) 67.0 ($\pm$2.7) 96.9 ($\pm$1.0) 68.9 ($\pm$2.7) 77.2 ($\pm$2.4) 24.9 ($\pm$2.5) 84.1 ($\pm$2.1) 27.1 ($\pm$2.6) 39.4 ($\pm$2.8)
Table 12: Paired 95% bootstrap confidence intervals for model differences. Each cell shows the 95% paired bootstrap CI on the first line and the one-sided bootstrap p-value below. MS = Mental State, BP = Behavior Prediction, JU = Behavior Judgment.
Model MS-BP BP-JU MS-JU BP-0.5 JU-0.5
DeepSeek-R1 [20.9, 25.9] (p << 0.001) [4.7, 11.3] (p << 0.001) [28.7, 34.3] (p << 0.001) [21.3, 26.4] (p=1.000) [13.0, 18.6] (p=1.000)
Meta-Llama-3.1-405B-Instruct-Turbo [36.8, 42.4] (p << 0.001) [45.2, 51.1] (p << 0.001) [85.9, 89.6] (p << 0.001) [5.4, 11.0] (p=1.000) [-41.7, -38.2] (p << 0.001)
Meta-Llama-3.1-8B-Instruct-Turbo [46.5, 52.7] (p << 0.001) [-20.0, -12.1] (p=1.000) [30.1, 37.1] (p << 0.001) [-14.3, -8.6] (p << 0.001) [1.7, 7.4] (p=0.999)
claude-3-5-sonnet-20240620 [28.2, 33.6] (p << 0.001) [38.9, 45.3] (p << 0.001) [70.4, 75.5] (p << 0.001) [14.3, 19.7] (p=1.000) [-27.6, -22.5] (p << 0.001)
claude-3-haiku-20240307 [60.7, 66.5] (p << 0.001) [3.6, 10.2] (p << 0.001) [67.4, 73.4] (p << 0.001) [-28.9, -23.9] (p << 0.001) [-35.4, -31.1] (p << 0.001)
claude-3-opus-20240229 [31.0, 36.6] (p << 0.001) [51.9, 57.9] (p << 0.001) [86.8, 90.4] (p << 0.001) [11.6, 17.1] (p=1.000) [-42.1, -38.7] (p << 0.001)
gpt-3.5-turbo-1106 [25.9, 32.0] (p << 0.001) [-24.3, -18.8] (p=1.000) [3.6, 11.2] (p << 0.001) [-43.9, -40.8] (p << 0.001) [-23.5, -18.2] (p << 0.001)
gpt-4-0125-preview [30.9, 36.4] (p << 0.001) [40.1, 47.0] (p << 0.001) [74.6, 79.5] (p << 0.001) [10.2, 15.7] (p=1.000) [-32.7, -28.1] (p << 0.001)
gpt-4.5-preview-2025-02-27 [26.6, 31.9] (p << 0.001) [38.0, 44.2] (p << 0.001) [67.7, 73.0] (p << 0.001) [15.1, 20.5] (p=1.000) [-25.9, -20.7] (p << 0.001)
gpt-4o-2024-05-13 [43.2, 49.1] (p << 0.001) [31.0, 37.5] (p << 0.001) [78.0, 82.7] (p << 0.001) [-3.4, 2.4] (p=0.366) [-36.8, -32.6] (p << 0.001)
gpt-4o-mini-2024-07-18 [49.3, 55.5] (p << 0.001) [14.7, 21.9] (p << 0.001) [67.9, 73.5] (p << 0.001) [-12.2, -6.5] (p << 0.001) [-30.0, -25.2] (p << 0.001)
gpt-5-2025-08-07 [31.4, 36.8] (p << 0.001) [21.3, 27.6] (p << 0.001) [55.7, 61.4] (p << 0.001) [11.7, 17.2] (p=1.000) [-12.9, -7.2] (p << 0.001)
o1-2024-12-17-high [35.7, 41.2] (p << 0.001) [24.0, 30.0] (p << 0.001) [62.6, 68.1] (p << 0.001) [7.4, 13.0] (p=1.000) [-19.4, -14.0] (p << 0.001)
o1-2024-12-17 [37.0, 42.6] (p << 0.001) [23.4, 29.2] (p << 0.001) [63.4, 68.9] (p << 0.001) [6.0, 11.6] (p=1.000) [-20.3, -14.8] (p << 0.001)
o1-mini-2024-09-12 [39.9, 46.0] (p << 0.001) [14.3, 21.1] (p << 0.001) [57.5, 63.9] (p << 0.001) [-8.1, -2.3] (p << 0.001) [-25.6, -20.4] (p << 0.001)
o1-preview-2024-09-12 [9.5, 13.4] (p << 0.001) [21.8, 27.6] (p << 0.001) [33.2, 39.0] (p << 0.001) [32.0, 36.2] (p=1.000) [6.6, 12.3] (p=1.000)
o3-mini-2025-01-31 [15.7, 21.5] (p << 0.001) [21.8, 29.0] (p << 0.001) [40.5, 47.5] (p << 0.001) [14.1, 19.6] (p=1.000) [-11.5, -5.8] (p << 0.001)
Q.3 Scenario-level accuracy with bootstrap-based error bars
To complement the aggregate analyses reported above in the main text, we further examine how these models behave across all scenario categories in SimpleToM, reported with error bars. Figure 1619 show per-scenario accuracies for mental state inference, behavior prediction, and judgment of behavior for four representative frontier models: Llama–3.1–405B, Claude 3.5 Sonnet, GPT–4o, and o1–preview.
Unlike Figure 4, which displayed raw accuracies only, the new scenario-level plots incorporate bootstrap 95% confidence intervals computed over the items within each scenario (using 5,0005{,}000 paired resamples). The error bars quantify uncertainty at the scenario level and reveal consistent patterns: (1) Mental state accuracy remains tightly concentrated near ceiling across all models and scenarios; (2) Behavior prediction accuracy varies substantially across scenarios, with larger confidence intervals indicating greater model instability; and (3) Judgment accuracy is uniformly the lowest and the error bars illustrate some uncertainty in models’ evaluations of others’ behavior.
Figure 16: Scenario-level accuracy for Llama–3.1–405B with 95% bootstrap error bars.
Figure 17: Scenario-level accuracy for Claude 3.5 Sonnet with 95% bootstrap error bars.
Figure 18: Scenario-level accuracy for GPT–4o with 95% bootstrap error bars.
Figure 19: Scenario-level accuracy for o1–preview with 95% bootstrap error bars.